Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Answer by fractalier(6550)

Let n be the number of nickels and d be the number of dimes.
Their values are then 5n and 10d.
The equations we write are
5n + 10d = 350 (thinking in cents now)
n + d = 50
Let's multiply the bottom one by 5 and subtract it from the top one...
5n + 10d = 350
-(5n + 5d = 250)
we get
5d = 100
d = 20 so then we get
n = 30